OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 5) and (–2, 11)?

OpenStudy (anonymous):

3x+y=5

OpenStudy (anonymous):

y – 11 = –3(x – 2) y – 5 = –3(x + 2) y – 5 = 3(x + 2) y – 11 = –3(x + 2)

OpenStudy (anonymous):

y – 11 = –3(x + 2)

OpenStudy (dumbcow):

slope: \[m = \frac{y2-y1}{x2-x1}\] point-slope form: \[y-y1 = m(x-x1)\]